### v2

# Einstein connection of nonsymmetric Riemannian manifold

Abstract
A. Einstein considered a nonsymmetric (0,2)-tensor $G=g+F$, where $g$ is a pseudo-Riemannian metric and $F\neq 0$ is skew-symmetric, and a linear connection $\nabla$ with torsion $T$ such that $(\nabla_{X}\,G)(Y,Z)=-G(T(X,Y),Z)$. M. Prvanović (1995) obtained the explicit form of the Einstein connection of an almost Hermitian manifold. In this paper, first, we present the result above in coordinate-free form, and then extend it to almost contact metric manifolds satisfying the $f^{2}$-torsion condition $T(f^{2}X,Y)=T(X,f^{2}Y)=f^{2}T(X,Y)$, where $F(X,Y)=g(X,fY)$. We then derive the Einstein connection of nonsymmetric pseudo-Riemannian, in particular, weak almost Hermitian manifolds $(M,f,g)$, satisfying the $f^{2}$-torsion condition, give explicit formulas for the torsion in terms of $\nabla^{g}F$, $dF$ and a new (1,1)-tensor $\widetilde{Q}=-f^{2}-{I}$, and show that in the almost Hermitian case, our results reduce to the coordinate-free form of Prvanović’s solution. Finally, we discuss special Einstein connections, and indicate the Gray-Hervella classes. Illustrative examples are given, including the construction of a weighted product.
Keywords: nonsymmetric pseudo-Riemannian manifold, weak almost Hermitian manifold, almost contact metric structure, Einstein connection, $f^{2}$-torsion condition.
MSC (2020) 53B05; 53C15; 53C21.
1 Introduction
A. Einstein worked on various versions of the Unified Field Theory (Nonsymmetric Gravitational Theory – NGT), see [2]. This theory was intended to unite the gravitation theory, to which General Relativity is related, and the theory of electromagnetism. Beginning in 1950, A. Einstein used a real nonsymmetric basic (0,2)-tensor $G$, whose symmetric part $g$ is associated with gravity, and the skew-symmetric part $F$ is associated with electromagnetism. In his attempt to construct a Unified Field Theory, (briefly NGT), A. Einstein [2] considered a snooth manifold $(M,G=g+F)$ with a linear connection $\nabla$ with torsion $T(X,Y)=\nabla_{X}Y-\nabla_{Y}X-[X,Y]$ satisfying the following condition:
$(\nabla_{X}G)(Y,Z)=-G(T(X,Y),Z)\quad(X,Y,Z\in\mathfrak{X}_{M}).$ (1.1)
called an Einstein connection. The idea of a nonsymmetric basic tensor was revisited by J.W. Moffat [7], allowing for richer gravitational dynamics. Results by T. Jansen and T. Prokopec [6] point to potential improvements and challenges within NGT. M.I. Wanas et al. [11] studied path equations of a particle in Einstein’s nonsymmetric geometry. Recent approaches to modified gravity often rely on differential geometry, including torsion and non-metricity, as a natural extension of General Relativity. Connections with totally skew-symmetric torsion, i.e., $T(X,Y,Z)$ is a 3-form, satisfying T(X,Y,Z)=T(X,Z,Y)=T(Y,X,Z)T(X,Y,Z)=-T(X,Z,Y)=-T(Y,X,Z), are important due to the relations with mathematical physics, as they admit a coupling to spinor fields and lead to holonomy and rigidity phenomena. In this case, the dimension of the space of parallel spinors coincides with the number of preserved supersymmetries. Such connections arise naturally in supersymmetric string theories and non-linear $\sigma$-models, as well as in gravitational models [4]. S. Ivanov and M. Lj. Zlatanović (see [5]) presented conditions for the existence and uniqueness of the Einstein connection with totally skew-symmetric torsion on a manifold $(M,G=g+F)$ and gave its explicit expression using an almost contact metric structure. In this case, the torsion is determined by $dF$: $T(X,Y,Z)=-\frac{1}{3}\,dF(X,Y,Z)$, and the difference tensor is given by
$K_{X}Y=\tfrac{1}{2}\big\{T({f}X,Y)-T(X,{f}Y)+T(X,Y)\big\}.$ (1.2)
Totally skew-symmetry of $K$ implies totally skew-symmetry of $T$, but the converse does not hold in general. In this paper, we do not assume that the torsion is totally skew-symmetric.
A linear connection $\nabla$} on a Riemannian manifold $(M,g)$ is said to be statistical if it is torsionless and tensor $({\nabla}g)(X,Y,Z):=({\nabla}_{X}\,g)(Y,Z)$ is symmetric in all its entries, equivalently, the cubic form $K(X,Y,Z):=g(K(X,Y),Z)$ (associated with the difference tensor (2.9) of $\nabla$) is symmetric. The statistical structure has applications in theory of probability and statistics as well as in information geometry. Unfortunately, we cannot directly apply the statistical structure to Einstein’s nonsymmetric geometry: if the cubic form $K(X,Y,Z)$ is symmetric, then by (2.11) and (2.12), $T=K=0$ is true; hence, $\nabla=\nabla^{g}$.
Weak metric structures, developed by V. Rovenski et. al, see [9], generalize the classical metric $f$-structures and are well suited for a study of $G=g+F$ in NGT with an arbitrary skew-symmetric tensor $F$. A weak almost Hermitian manifold $(M,f,g)$ is a Riemannian manifold $(M,g)$ of dimension $2m\geq 4$ endowed with non-singular skew-symmetric endomorphism $f$ and the 2-form $F$ such that $F(X,Y)=g(X,fY)$, see [12]. V. Rovenski and M. Zlatanović [12, 10] were the first to apply the weak $f$-structure, in particular, the weak almost contact structure, to NGT of totally skew-symmetric torsion with $F(X,Y)=g(X,fY)$.
M. Prvanović [8] obtained the explicit form of the Einstein connection of almost Hermitian manifold, and using its torsion tensor, characterized the sixteen classes (by A. Gray and L.M. Hervella [3]) of almost Hermitian manifold.
In this paper, we obtain the explicit form of the Einstein connection of an NGT space, in particular, a weak almost Hermitian manifold, satisfying some natural condition, called the $f^{2}$-torsion condition. This condition disappears for the classical case, hence our result generalizes the result by M. Prvanović. We also present the explicit form of the Einstein connection of an almost contact metric manifold satisfying the $f^{2}$-torsion condition.
The paper has an Introduction and six sections. In Section 2 we review basics of NGT, give examples, prove several lemmas, and discuss the relation between torsion and the contorsion, $K$, of an Einstein connection emphasizing the role of $f^{2}$-torsion condition. In Section 3 we deduce equations for Einstein connection of an almost contact metric manifold with the $f^{2}$-torsion condition (Theorem 3.3). In Section 4 we generalize the result of M. Prvanović and deduce equations for an Einstein connection with the $f^{2}$-torsion condition of nonsymmetric pseudo-Riemannian manifold (Theorem 4.3), in particular, for weak almost Hermitian manifold (Corollary 4.5). Long expressions of Section 4 are given in Section 5.
2 Einstein’s nonsymmetric geometry
A manifold $(M,G=g+F)$ equipped with an Einstein connection $\nabla$ different from the Levi-Civita connection $\nabla^{g}$ is called an NGT space. The basic (0,2)-tensor $G$ of an NGT space $(M,G=g+F)$ decomposes into two parts, the symmetric part $g$ (pseudo-Riemannian metric) and the skew-symmetric part $F$ (fundamental 2-form), that is,
$g(X,Y)=\tfrac{1}{2}\big[G(X,Y)+G(Y,X)\big],\quad F(X,Y)=\tfrac{1}{2}\big[G(X,Y)-G(Y,X)\big].$
We assume that the skew-symmetric part, $F$ is non-zero, may have arbitrary (not necessarily constant) rank, and in particular, may be non-degenerate (that is, of maximal rank). Therefore, we obtain a (1,1)-tensor ${f}\neq 0$ (of not necessarily constant rank) determined by
$g(X,fY)=F(X,Y)\quad(X,Y\in\mathfrak{X}_{M}).$
Since $F$ is skew-symmetric, the tensor f{f} is also skew-symmetric: g(fX,Y)=g(X,fY)g({f}X,Y)=-g(X,{f}Y). Note that f=Af=-A, where the tensor $A$ is given in [5] by the equality $g(AX,Y)=F(X,Y)$.
The torsion of a linear connection $\nabla$ on $M$ is given by
$T(X,Y)=\nabla_{X}Y-\nabla_{Y}X-[X,Y].$
The covariant derivative an $(p,k)$-tensor $S$ for $p=0,1$ is an $(p,k+1)$-tensor $\nabla S$:
$(\nabla S)(Y,X_{1},\ldots,X_{k})=\nabla_{Y}(S(X_{1},\ldots,X_{k}))-\sum\nolimits_{\,i=1}^{k}S(X_{1},\ldots,\nabla_{Y}X_{i},\ldots,X_{k}).$ (2.1)
Separating symmetric and skew-symmetric parts of (1.1), we express the covariant derivatives $\nabla g$ and $\nabla F$ in terms of the (0,3)-torsion tensor $T(X,Y,Z)=g(T(X,Y),Z)$:
$2\,(\nabla_{X}\,g)(Y,Z)$ $=T(Z,X,\,Y+fY)-T(X,Y,\,Z+fZ),$ (2.2)
$2\,(\nabla_{Z}\,F)(X,Y)$ $=-T(Z,X,\,Y+fY)-T(Y,Z,\,X+fX).$ (2.3)
Recall the co-boundary formula for a 2-form $F$ (without the coefficient 3, unlike [1]):
$dF(X,Y,Z)$ $=X(F(Y,Z))+Y(F(Z,X))+Z(F(X,Y))$
$\ -F([X,Y],Z)-F([Z,X],Y)-F([Y,Z],X).$ (2.4)
The equality (2) yields
$dF(X,Y,Z)=(\nabla^{g}_{X}\,F)(Y,Z)+(\nabla^{g}_{Y}\,F)(Z,X)+(\nabla^{g}_{Z}\,F)(X,Y).$ (2.5)
The metricity condition (1.1) can be writen in the form
$(\nabla_{X}(g+F))(Y,Z)=-\,T(X,Y,Z)-T(X,Y,fZ).$ (2.6)
Taking the cyclic sum in (2.6) and applying the equality
$dF(X,Y,Z)$ $=\,T(X,Y,fZ)+\,T(Y,Z,fX)+\,T(Z,X,fY)$
$+(\nabla_{X}F)(Y,Z)+(\nabla_{Y}F)(Z,X)+(\nabla_{Z}F)(X,Y),$
which follows from (2.3) and (2.5), we obtain
$\quad(\nabla_{X}g)(Y,Z)+(\nabla_{Y}g)(Z,X)+(\nabla_{Z}g)(X,Y)$ (2.7)
$=-\,dF(X,Y,Z)-\,T(X,Y,Z)-\,T(Y,Z,X)-\,T(Z,X,Y).$
Since the left hand side of (2.7) is symmetric, while the right hand side is skew-symmetric, we get the following two relations, see [5, Eq. (3.3)]:
$(\nabla_{X}\,g)(Y,Z)+(\nabla_{Y}\,g)(Z,X)+(\nabla_{Z}\,g)(X,Y)=0,$
and
$dF(X,Y,Z)=-T(X,Y,Z)-T(Y,Z,X)-T(Z,X,Y).$ (2.8)
The difference tensor $K$ of a linear connection $\nabla$ and the Levi-Civita connection $\nabla^{g}$ is
$K_{X}Y:=\nabla_{X}Y-\nabla^{g}_{X}Y.$ (2.9)
The Einstein connection $\nabla$ is represented in [5, Eq. (3.7)] using the torsion $T$ as
$g(\nabla_{X}Y,Z)$ $=g(\nabla^{g}_{X}Y,Z)+\tfrac{1}{2}\big\{T(X,Y,Z)-T(Z,X,{f}Y)+T(Y,Z,{f}X)\big\}.$ (2.10)
By (2.10), the difference (0,3)-tensor $K(X,Y,Z)=g(K_{X}Y,Z)$ and the torsion tensor of an Einstein connection $\nabla$ are expressed linearly in terms of each other:
$2\,K(X,Y,Z)$ $=T(X,Y,Z)-T(Z,X,{f}Y)+T(Y,Z,{f}X),$ (2.11)
$T(X,Y,Z)$ $=K(X,Y,Z)-K(Y,X,Z)\quad\Longleftrightarrow\quad T(X,Y)=K_{X}Y-K_{Y}X.$ (2.12)
Using various presentations of tensors and (2.1), we obtain
$(\nabla^{g}F)(Z,X,Y)=(\nabla^{g}_{Z}\,F)(X,Y)=g(X,(\nabla^{g}_{Z}\,f)Y).$ (2.13)
Lemma 2.1.
For an Einstein connection $\nabla$, the following conditions are equivalent:
(i) K(X,Y,Z)=K(X,Z,Y)K(X,Y,Z)=-K(X,Z,Y),
(ii) $\nabla g=0$,
(iii) $(\nabla_{X}F)(Y,Z)=-(\nabla_{Y}F)(X,Z)$, or $g((\nabla_{X}f)Z,Y)=-g((\nabla_{Y}f)Z,X)$.
Proof.
We calculate using (2.1) and (2.9),
$(\nabla_{X}\,g)(Y,Z)$ $=X(g(Y,Z))-g(\nabla^{g}_{X}Y+K_{X}Y,Z)-g(Y,\nabla^{g}_{X}Z+K_{X}Z)$
$=(\nabla^{g}_{X}\,g)(Y,Z)-g(K_{X}Y,Z)-g(K_{X}Z,Y).$
By this and $\nabla^{g}\,g=0$, the conditions (i) and (ii) are equivalent. The equations (2.2) and (2.3) yield, see [5, Eq. (3.4)],
$(\nabla_{Z}\,g)(X,Y)=(\nabla_{X}F)(Y,Z)+(\nabla_{Y}F)(X,Z).$ (2.14)
By (2.14) and (2.13), the conditions (ii) and (iii) are equivalent. ∎
Remark $2.$
When $\nabla$ preserves the metric tensor: $\nabla g=0$, they call $K$ the contorsion tensor; in this case the tensors $K$ and $T$ are related by
$2\,K(Y,Z,X)=T(X,Y,Z)+T(Y,Z,X)-T(Z,X,Y).$ (2.15)
Comparing (2.15) and (2.11) yields the following identity for the torsion of an Einstein connection satisfying $\nabla g=0$:
$T(Z,X,Y)-T(Y,Z,X)=T(Y,Z,{f}X)-T(Z,X,{f}Y).$ (2.16)
Example 2.3.
Let $(M,f,g)$ be a weak nearly Kähler manifold, considered as a nonsymmetric Riemannian manifold $(M,G=g+F)$ with $F(X,Y)=g(X,fY)$. If $\nabla$ is an Einstein connection on $M$ with totally skew-symmetric torsion $T$, then the $f$-torsion condition holds:
$T(fX,Y)=T(X,fY)=-f\,T(X,Y),$ (2.17)
and the torsion of $\nabla$ is determined by $T(X,Y,Z)=-\tfrac{1}{3}\,dF(X,Y,Z)$, see Example 3.4 in [12]. In this case, (1.2) reduces to $K_{X}Y=\tfrac{1}{2}\,T(X,Y)$, and $\nabla_{X}Y=\nabla^{g}_{X}Y+\tfrac{1}{2}\,T(X,Y)$ holds.
Remark 2.4.
Since for an Einstein connection on $(M,G=g+F)$, the tensor $K(X,Y,Z)$ is not totally skew-symmetric, it is interesting to study its particular symmetries. The skew-symmetry of $K_{X}$ i.e., K(X,Y,Z)=K(X,Z,Y)K(X,Y,Z)=-K(X,Z,Y), see Lemma 2.1(i), by (2.11), reduces to
$T(X,Y,Z)-T(Z,X,Y)-T(Z,X,fY)+T(X,Y,fZ)=0.$ (2.18)
Using (2.18) in (2.3), we obtain
$(\nabla_{X}F)(Y,Z)=-T(X,Y,Z)-T(X,Y,fZ).$
If (2.18) is true, then $K_{X},K_{Y}]:TM\to TM$ is a skew-symmetric endomorphism:
$[K_{X},K_{Y}](Z,Z)=0,\quad X,Y,Z\in\mathfrak{X}_{M}.$ (2.19)
If a nonsymmetric Riemannian manifold $(M,G=g+F)$ admits an Einstein connection $\nabla$ with totally skew-symmetric torsion, and (2.17) holds, then (2.19) is true. Indeed, using (2.17) and the totally skew-symmetric torsion property in the LHS of (2.18) gives zero.
The following skew-symmetry of the difference tensor $K$ is obtained using (2.12):
$K_{X}Y=-K_{Y}X,\quad\text{or, equivalently,}\quad K_{X}Y=\tfrac{1}{2}\,T(X,Y),$ (2.20)
and characterizes special Einstein connections, see [8]. This condition expresses the skew–symmetry of $K$ only in the first two arguments, and doesn’t imply the total skew–symmetry of $K(X,Y,Z)$, which requires, in addition, condition (2.18). In this case, torsion tensor becomes totally skew–symmetric and satisfies the $f$-torsion condition, see Lemma 2.5.
Lemma 2.5.
If the tensor $K$ is totally skew-symmetric, then the torsion tensor $T$ is totally skew-symmetric, and the $f$-torsion condition (2.17) is true.
Proof.
Let the tensor $K(X,Y,Z)$ be totally skew-symmetric. Then we obtain
$K(X,Y,Z)+K(Y,X,Z)=0\implies T(Y,Z,fX)-T(Z,X,fY)=0.$
By Lemma 2.1, we have $\nabla g=0$. Combining this identity with (2.16) and using the assumption that $K(X,Y,Z)$ is totally skew-symmetric, we deduce
$T(X,Y,Z)=-T(X,Z,Y)\quad\text{and}\quad T(X,Y,fZ)=-T(X,Z,fY).$
So, $T(X,Y,Z)$ is totally skew-symmetric and the $f$-torsion condition (2.17) is true. ∎
Proposition 2.6.
Let the condition (2.18) be true. Then we get the equality
$K(X,Y,Z)=T(Z,Y,X)-\tfrac{1}{2}\,dF(X,Y,Z);$ (2.21)
therefore, the tensor $K$ is totally skew-symmetric if and only if the torsion $T$ is totally skew-symmetric. Moreover, the total skew-symmetry of $T$ implies the condition (2.17).
Proof.
Applying (2.12) to (2.8) gives
$K(X,Y,Z)-K(Y,X,Z)+K(Y,Z,X)-K(Z,Y,X)$
$+$ $K(Z,X,Y)-K(X,Z,Y)=-dF(X,Y,Z).$ (2.22)
Then applying (2.18) to (2), yields
$K(X,Y,Z)-K(Y,X,Z)-K(Y,X,Z)+K(Z,X,Y)+K(Z,X,Y)+K(X,Y,Z)$
$=2\big(K(X,Y,Z)-K(Y,X,Z)+K(Z,X,Y)\big)=-dF(X,Y,Z).$ (2.23)
Using (2.12) to (2) gives (2.21). By the above, the total skew-symmetry of $T$ implies the total skew-symmetry of $K$. By Lemma 2.5, the $f$-torsion condition (2.17) is true. ∎
Proposition 2.7 (see Proposition 2 of [8]).
For an Einstein connection $\nabla$ with torsion $T$, we have
$2(\nabla^{g}_{X}F)(Y,Z)=$ $-T(Z,X,Y)-T(X,Y,Z)-T(fZ,X,fY)$
$-T(X,fY,fZ)+T(Y,fZ,fX)+T(fY,Z,fX).$ (2.24)
Proof.
By (2.1), we have
$(\nabla_{X}F)(Y,Z)=X\big(F(Y,Z)\big)-F(\nabla_{X}Y,Z)-F(Y,\nabla_{X}Z).$
Using the relation (2.9), we obtain
$(\nabla_{X}F)(Y,Z)$ $=X\big(F(Y,Z)\big)-F(\nabla^{g}_{X}Y+K_{X}Y,Z)-F(Y,\nabla^{g}_{X}Z+K_{X}Z)$
$=(\nabla^{g}_{X}F)(Y,Z)-F\big(K_{X}Y,Z\big)-F\big(Y,K_{X}Z\big)$
$=(\nabla^{g}_{X}F)(Y,Z)-g\big(K_{X}Y,fZ\big)+g\big(fY,K_{X}Z\big)$
$=(\nabla^{g}_{X}F)(Y,Z)-K(X,Y,fZ)+K(X,Z,fY),$ (2.25)
Finally, (2.7) follows from (2.3), (2.11) and (2). ∎
We represent the solution (an explicit form of the Einstein connection of an almost Hermitian manifold) of M. Prvanović [8] in the coordinate-free form.
Theorem 2.8 (see Theorem 1 of [8]).
Let $(M^{2n},J,g)$ be an almost Hermitian manifold considered as a nonsymmetric Riemannian manifold $(M,G=g+F)$, where $F(X,Y)=g(X,JY)$ Then an Einstein connection $\nabla$ on $M$ is given by
$2\,T(Y,Z,X)=$ $\ 2\,(\nabla^{g}_{JX}F)(JY,Z)-(\nabla^{g}_{JY}F)(JZ,X)-(\nabla^{g}_{JZ}F)(X,JY)$
$-(\nabla^{g}_{Y}F)(Z,X)-(\nabla^{g}_{Z}F)(X,Y).$ (2.26)
In view of (2.8), the Einstein connection $\nabla$ is special, see definition (2.20), if and only if the following identity holds:
$(\nabla^{g}_{X}F)(Y,Z)+(\nabla^{g}_{Y}F)(X,Z)=(\nabla^{g}_{JX}F)(JY,Z)+(\nabla^{g}_{JY}F)(JX,Z).$ (2.27)
Corollary 2.9 (see Theorem 2 of [8]).
Let $(M^{2n},J,g)$ be an almost Hermitian manifold considered as a generalized Riemannian manifold $(M,G=g+F)$, where $F(X,Y)=g(X,JY)$. Then the torsion of a special Einstein connection $\nabla$ on $M$, see (2.20), is given by
$2\,T(X,Y,Z)=(\nabla^{g}_{X}F)(Y,Z)-(\nabla^{g}_{JZ}F)(JX,Y)-(\nabla^{g}_{JY}F)(JX,Z).$
Among the sixteen Gray-Hervella classes [3] of almost Hermitian manifolds, the condition (2.27) of a special Einstein connection is satisfied for the following classes:
${\mathcal{W}}_{1},\ \ {\mathcal{W}}_{3},\ \ {\mathcal{W}}_{4},\ \ {\mathcal{W}}_{3}\oplus{\mathcal{W}}_{4},\ \ {\mathcal{W}}_{1}\oplus{\mathcal{W}}_{3},\ \ {\mathcal{W}}_{1}\oplus{\mathcal{W}}_{4},\ \ {\mathcal{W}}_{1}\oplus{\mathcal{W}}_{3}\oplus{\mathcal{W}}_{4};$
consequently, the Einstein connection on an almost Hermitian manifold belonging to any of these classes is a special Einstein connection, see Theorem 3 of [8].
3 Einstein connection of an almost contact metric manifold
Contact Riemannian geometry is a significant field in both pure mathematics and theoretical physics, for example, [1]. Developing the approach of [8], we deduce equations for Einstein connection of an almost contact metric manifold $(M^{2n+1},f,\xi,\eta,g)$. Here, $g$ is a Riemannian metric, $f$ is a skew-symmetric (1,1)-tensor of rank 2n2\,n, $\xi$ is a unit (Reeb) vector field, and $\eta$ is a 1-form such that $\eta(\xi)=1$. These tensors satisfy the relations
$f^{2}=-{I}+\eta\otimes\xi,\quad g(fX,fY)=g(X,Y)-\eta(X)\eta(Y),$ (3.1)
where I{I} is the identity map. Differentiating by $X$ the first identity of (3.1), yields
$\nabla^{g}_{X}f^{2}=\nabla^{g}_{X}(-{I}+\eta\otimes\xi)=\nabla^{g}_{X}(\eta\otimes\xi).$
Using the Leibniz rule, we have
$\nabla^{g}_{X}(\eta\otimes\xi)=(\nabla^{g}_{X}\eta)\otimes\xi+\eta\otimes\nabla^{g}_{X}\xi,\qquad\nabla^{g}_{X}f^{2}=(\nabla^{g}_{X}f)\,f+f\,\nabla^{g}_{X}f.$
Applying the equalities above to $Y$, yields
$(\nabla^{g}_{X}f)\,fY+f\,(\nabla^{g}_{X}f)Y=(\nabla^{g}_{X}\eta)(Y)\,\xi+\eta(Y)\,\nabla^{g}_{X}\xi.$ (3.2)
Definition 3.1.
A linear connection $\nabla$ satisfies the $f^{2}$-torsion condition if the following is true:
$T(f^{2}X,Y)=T(X,f^{2}Y)=f^{2}T(X,Y),$ (3.3)
which can be equivalently written as
$T(f^{2}X,Y,Z)=T(X,f^{2}Y,Z)=T(X,Y,f^{2}Z).$
If (3.3) is true for an Einstein connection, then, using (2.8) and (2.7), we have
$dF(f^{2}X,Y,Z)=dF(X,f^{2}Y,Z)=dF(X,Y,f^{2}Z),$
$(\nabla^{g}F)(f^{2}X,Y,Z)=(\nabla^{g}F)(X,f^{2}Y,Z)=(\nabla^{g}F)(X,Y,f^{2}Z).$
If the $f$-torsion condition (2.17) holds, then the $f^{2}$-torsion condition (3.3) also holds. The converse is not true in general.
Lemma 3.2.
The following conditions are true for all vector fields $X,Y,Z$ on an almost contact metric manifold:
$(\nabla^{g}_{X}F)(fY,fZ)$ $=-(\nabla^{g}_{X}F)(Y,Z)+\eta(Y)\,(\nabla^{g}_{X}F)(\xi,Z)+\eta(Z)\,(\nabla^{g}_{X}F)(Y,\xi),$ (3.4)
$(\nabla^{g}_{X}F)(fY,Z)$ $=(\nabla^{g}_{X}F)(Y,fZ)+\eta(Z)\,(\nabla^{g}_{X}F)(fY,\xi)-\eta(Y)\,(\nabla^{g}_{X}F)(\xi,fZ).$ (3.5)
Proof.
Using $\nabla^{g}g=0$ and $F(Y,Z)=g(Y,fZ)$, we have $(\nabla^{g}_{X}F)(Y,Z)=g\big(Y,(\nabla^{g}_{X}f)Z\big)$. Hence
$(\nabla^{g}_{X}F)(fY,fZ)=g\big(fY,(\nabla^{g}_{X}f)\,fZ\big).$
Substituting (3.2), we get
$(\nabla^{g}_{X}f)\,fZ=-f\,(\nabla^{g}_{X}f)Z+(\nabla^{g}_{X}\eta)(Z)\xi+\eta(Z)\nabla^{g}_{X}\xi.$
Using $g(\xi,fZ)=0$ gives
$(\nabla^{g}_{X}F)(fY,fZ)$ $=-\,g\left(fY,f\,(\nabla^{g}_{X}f)Z\right)+\eta(Z)\,g(fY,\nabla^{g}_{X}\xi).$
By $g(fX,fY)=g(X,Y)-\eta(X)\eta(Y)$ and $\eta(fZ)=0$, we obtain
$g\!\left(fY,f(\nabla^{g}_{X}f)Z\right)=g\big(Y,(\nabla^{g}_{X}f)Z\big)-\eta(Y)\eta\big((\nabla^{g}_{X}f)Z\big).$
Therefore,
$(\nabla^{g}_{X}F)(fY,fZ)=-g\big(Y,\nabla^{g}_{X}f)Z\big)+\eta(Y)\eta\big((\nabla^{g}_{X}f)Z\big)+\eta(Z)\,g(fY,\nabla^{g}_{X}\xi).$ (3.6)
Finally, substituting the equalities
$g\big(Y,(\nabla^{g}_{X}f)Z\big)$ $=(\nabla^{g}_{X}F)(Y,Z),\quad\eta\big((\nabla^{g}_{X}f)Z\big)=g\big(\xi,(\nabla^{g}_{X}f)Z\big)=(\nabla^{g}_{X}F)(\xi,Z),$
$(\nabla^{g}_{X}F)(Y,\xi)$ $=g\big(Y,(\nabla^{g}_{X}f)\xi\big)=-g\big(Y,f\nabla^{g}_{X}\xi\big)=g(\nabla^{g}_{X}\xi,fY).$
in (3.6), yields (3.4). Taking $Z\to fZ$ in (3.4), we get (3.5). ∎
Theorem 3.3.
Let $(M^{2n+1},f,\xi,\eta,g)$ be an almost contact metric manifold, considered as an NGT space $(M,G=g+F,\nabla)$ with $F(X,Y)=g(X,fY)$. If the Einstein connection $\nabla$ satisfies the $f^{2}$-torsion condition (3.3), then
$\nabla^{g}\,\xi=\nabla^{g}\,\eta=0,\quad T(\xi,\cdot,\cdot)=T(\cdot,\cdot,\xi)=0,\quad(\nabla^{g}F)(\xi,\cdot,\cdot)=(\nabla^{g}F)(\cdot,\cdot,\xi)=0.$ (3.7)
In particular, the torsion $T$ is horizontal, and for $X,Y,Z\perp\xi$ it is given by the same formula (2.8) as for the almost Hermitian case:
$2\,T(Y,Z,X)=$ $\ 2\,(\nabla^{g}_{fX}F)(fY,Z)-(\nabla^{g}_{fY}F)(fZ,X)-(\nabla^{g}_{fZ}F)(X,fY)$
$-(\nabla^{g}_{Y}F)(Z,X)-(\nabla^{g}_{Z}F)(X,Y).$ (3.8)
Proof.
We write $Z=f^{2}Z^{\prime}+\eta(Z)\xi$, where $Z^{\prime}\perp\xi$, and by (3.3), we get
$T(\xi,Z,\xi)=T(\xi,f^{2}Z^{\prime},\xi)+\eta(Z)\,T(\xi,\xi,\xi)=T(\xi,f^{2}Z^{\prime},\xi)=T(f^{2}\xi,Z^{\prime},\xi)=0.$ (3.9)
Using the $f^{2}$-torsion condition (3.3) again and (3.9), we obtain
$T(X,Z,\xi)$ $=T(X,f^{2}Z^{\prime},\xi)+\eta(Z)\,T(X,\xi,\xi)=T(f^{2}X,Z^{\prime},\xi)=0,$
$T(\xi,Z,X)$ $=T(\xi,f^{2}Z^{\prime},X)+\eta(Z)\,T(\xi,\xi,X)=T(f^{2}\xi,Z^{\prime},X)=0.$
Hence,
$T(\xi,Z,\xi)=T(X,Z,\xi)=T(\xi,Z,X)=0.$ (3.10)
Substituting $X=\xi$ into (2.7), we obtain
$\nabla^{g}_{\xi}\,F=0,\quad\mbox{that is,}\quad\nabla^{g}_{\xi}\,f=0.$ (3.11)
Therefore, the $\xi$-trajectories are $g$-geodesics. From (2.7), we have
$(\nabla^{g}_{X}F)(Y,Z)-(\nabla^{g}_{X}F)(fY,fZ)=T(Y,fZ,fX)+T(fY,Z,fX).$ (3.12)
Taking into account (3.4), (3.10) and (3.11), from (3.12) we obtain
$2(\nabla^{g}_{X}F)(Y,Z)=T(Y,fZ,fX)+T(fY,Z,fX).$ (3.13)
Replacing $X\to fX$ and $Y\to fY$ in (3.13), we get
$2(\nabla^{g}_{fX}F)(fY,Z)=-T(fY,fZ,X)+T(Y,Z,X)=-2(\nabla^{g}_{fX}F)(fZ,Y).$
We calculate, following the approach of [8]:
$(\nabla^{g}_{X}F)(fY,fZ)=-(\nabla^{g}_{X}F)(Y,Z),\qquad(\nabla^{g}_{X}F)(fY,Z)=(\nabla^{g}_{X}F)(Y,fZ).$ (3.14)
From (1.1) with $G=g+F$ we obtain
$2\,(\nabla^{g}_{X}F)(Y,Z)=$ $\,-T(Z,X,Y)-T(X,Y,Z)-T(X,fY,fZ)$
$-T(fZ,X,fY)+T(Y,fZ,fX)+T(fY,Z,fX).$ (3.15)
Using (3.15), we get
$(\nabla^{g}_{X}F)(Y,Z)-(\nabla^{g}_{X}F)(fY,fZ)=T(Y,fZ,fX)+T(fY,Z,fX).$
Substituting this into (3.15) and using (3.14), we obtain
$T(X,Y,Z)+T(Z,X,Y)+T(X,fY,fZ)+T(fZ,X,fY)=0.$
Thus (3.15) reduces to
$2\,(\nabla^{g}_{X}F)(Y,Z)=T(Y,fZ,fX)+T(fY,Z,fX).$ (3.16)
The cyclic sum of (3.16) is
$(\nabla^{g}_{X}F)(Y,Z)+(\nabla^{g}_{Y}F)(Z,X)+(\nabla^{g}_{Z}F)(X,Y)$
$=-T(X,Y,Z)-T(Y,Z,X)-T(Z,X,Y).$ (3.17)
Replacing $Y\to fY$ and $Z\to fZ$ in (3) gives
$(\nabla^{g}_{X}F)(fY,fZ)+(\nabla^{g}_{fY}F)(fZ,X)+(\nabla^{g}_{fZ}F)(X,fY)$
$=T(X,Y,Z)+T(Z,X,Y)-T(fY,fZ,X).$ (3.18)
Replacing $X\to fX$ and $Y\to fY$ in (3.16), we obtain
$T(fY,fZ,X)=T(Y,Z,X)-2\,(\nabla^{g}_{fX}F)(fY,Z).$ (3.19)
Substituting (3.19) into (3), gives the formula (3.3). ∎
Remark 3.4.
Although the $f^{2}$-torsion condition becomes trivial for almost Hermitian manifolds, it is strong enough for almost contact metric manifolds. In Theorem 3.3, our manifold is locally the direct product of $\mathbb{R}$ and an almost Hermitian manifold $(\bar{M}^{2n},J,\bar{g})$, and $\nabla$, being an Einstein connection on the $2n$-dimensional factor, is given by (2.8), where $J$ is replaced by $f$. Consequently, the 1-form $\eta$ is closed, $d\eta=0$. The distribution $\ker\eta$ is involutive; thus, it is Frobenius integrable. The integral submanifold $\bar{M}^{2n}$ is almost Hermitian.
We work with an Einstein connection satisfying the $f^{2}$-torsion condition, without assuming any symmetry of the torsion tensor. If, in addition, the torsion $T$ is totally skew–symmetric, then the general formula (3.3) reduces to (3.21), and $\nabla^{g}_{\xi}F=0$. This additional assumption leads to the almost–nearly cosymplectic case, which we recall below from [5].
Corollary 3.5.
Under the conditions of Theorem 3.3, the torsion $T$ is totally skew-symmetric if and only if (3.7) is true, and the following Codazzi-type condition holds:
$(\nabla^{g}_{X}F)(Y,Z)=(\nabla^{g}_{Y}F)(Z,X)=(\nabla^{g}_{Z}F)(X,Y)=\tfrac{1}{3}\,dF(X,Y,Z).$ (3.20)
In this case, the torsion is given by
$T(X,Y,Z)=-(\nabla^{g}_{Z}F)(X,Y)=-\tfrac{1}{3}\,dF(X,Y,Z).$ (3.21)
Proof.
By Theorem 3.3, we have $\nabla^{g}_{\xi}\,F=0$ and $\nabla^{g}\,\xi=0$. Let’s consider the $f^{2}$-torsion condition under the additional assumption that $T$ is totally skew-symmetric. Applying T(Y,X,Z)=T(Y,Z,X)T(Y,X,Z)=-T(Y,Z,X), from (3.3) we have
$0=3(\nabla^{g}_{fX}F)(fY,Z)+3(\nabla^{g}_{fZ}F)(fY,X)-(\nabla^{g}_{Z}F)(X,Y)-(\nabla^{g}_{X}F)(Z,Y).$ (3.22)
Taking $X\to fX$ and $Y\to fY$ in (3.22) and using (3.7) and the equality $(\nabla^{g}_{X}F)(fY,fZ)=-(\nabla^{g}_{X}F)(Y,Z)$, see (3.4), we get
$0=\ 3\,(\nabla^{g}_{X}F)(Y,Z)-3\,(\nabla^{g}_{fZ}F)(Y,fX)+(\nabla^{g}_{Z}F)(X,Y)-(\nabla^{g}_{fX}F)(Z,fY).$
Adding the equation above to (3.22), gives
$(\nabla^{g}_{fX}F)(fY,Z)=-(\nabla^{g}_{X}F)(Y,Z).$ (3.23)
Substituting (3.23) into (3.22), we obtain $(\nabla^{g}_{X}F)(Y,Z)=(\nabla^{g}_{Z}F)(X,Y)$; hence, in view of (2.5), the condition (3.20) is true. Finally, from (3.3) and (3.23) we get $T(Y,Z,X)=-(\nabla^{g}_{X}F)(Y,Z)$ that, in view of (3.20), completes the proof of (3.21). ∎
An almost contact metric manifold $(M^{2n+1},f,\xi,\eta,g)$ is said to be almost-nearly cosymplectic, see [5], if the Levi–Civita covariant derivative of $f$ satisfies
$g\bigl((\nabla^{g}_{X}f)Y,Z\bigr)=\tfrac{1}{3}\,dF(fX,fY,Z)+\tfrac{1}{6}\,\eta(Z)\,d\eta(Y,fX)-\tfrac{1}{2}\,\eta(Y)\,d\eta(fZ,X),$
where $F(X,Y)=g(X,fY)$, in particular, $\xi$ is a Killing vector field.
Corollary 3.6 (see [5]).
Let $(M^{2n+1},f,\xi,\eta,g)$ be an almost contact metric manifold with fundamental 22–form $F$, considered as a nonsymmetric Riemannian manifold $(M,G=g+F)$. Then an Einstein connection $\nabla$ on $M$ has a totally skew-symmetric torsion $T$ if and only if the manifold is almost–nearly cosymplectic, in this case, $T(X,Y,Z)=-\tfrac{1}{3}\,dF(X,Y,Z)$.
The next consequences of Theorem 3.3 explicitly present the special Einstein connection with the totally skew-symmetric torsion satisfying the $f^{2}$-torsion condition.
Corollary 3.7.
Let $\nabla$ be a special Einstein connection satisfying the $f^{2}$-torsion condition (3.3) on an almost contact metric manifold $(M^{2n+1},f,\xi,\eta,g)$, considered as an NGT space $(M,G=g+F,\nabla)$, where $F(X,Y)=g(X,fY)$. Then (3.7) is true, and the torsion $T$ of $\nabla$ is given by
$2\,T(X,Y,Z)=(\nabla^{g}_{X}F)(Y,Z)-(\nabla^{g}_{fZ}F)(fX,Y)-(\nabla^{g}_{fY}F)(fX,Z).$ (3.24)
Proof.
By Theorem 3.3, we have $\nabla^{g}_{\xi}\,F=0$ and $\nabla^{g}\,\xi=0$. The rest of the proof is the same as the proof of Corollary 2.9. ∎
Corollary 3.8.
Under the conditions of Theorem 3.3, let $\nabla$ be a special Einstein connection. Then the torsion $T$ is totally skew-symmetric if and only if (3.7) is true; in this case,
$T(X,Y,Z)=\tfrac{1}{2}(\nabla^{g}_{X}F)(Y,Z)=\tfrac{1}{6}\,dF(X,Y,X).$ (3.25)
Proof.
Assume that the torsion $T$ of $\nabla$ is totally skew-symmetric. Applying T(X,Y,Z)=T(X,Z,Y)T(X,Y,Z)=-T(X,Z,Y) to (3.24), we obtain
$(\nabla^{g}_{fZ}F)(fX,Y)+(\nabla^{g}_{fY}F)(fX,Z)=0.$ (3.26)
Substituting the identity (3.26) into (3.24) yields the first equality of (3.25). Taking $Z\to fZ$ and $Y\to fY$ in (3.26) and using the equalities (3.7), $(\nabla^{g}_{X}F)(fY,fZ)=-(\nabla^{g}_{X}F)(Y,Z)$, see (3.4), and $\nabla^{g}_{\xi}\,F=\nabla^{g}_{X}F(\xi,\,\cdot)=0$, in view of (2.5), we get the second equality of (3.25):
$(\nabla^{g}_{X}F)(Y,Z)=(\nabla^{g}_{Y}F)(Z,X)=(\nabla^{g}_{Z}F)(X,Y)=\tfrac{1}{3}\,dF(X,Y,Z).$
Conversely, if (3.25) holds, then $T$ is totally skew-symmetric, since $(\nabla^{g}_{X}F)(Y,Z)$ is skew-symmetric in $Y,Z$. The conditions $\nabla^{g}_{\xi}\,F=0$ and $\nabla^{g}\,\xi=0$ follow from Theorem 3.3. ∎
4 Einstein connection of an NGT space
In this section, we present explicitly an Einstein connection of a nonsymmetric pseudo-Riemannian space $(M,G=g+F)$, in particular, of a weak almost Hermitian manifold.
A pseudo-Riemannian manifold $(M^{2n},g)$ endowed with a skew-symmetric (1,1)-tensor $f$ of maximal rank $2n$ ($f^{2}$ is not necessarily equal to I-{I}) is called a weak almost Hermitian manifold. If $\nabla^{g}J=0$, then such $(M^{2n},J,g)$ is called a weak Kähler manifold, see [9].
The $f^{2}$-torsion condition (3.3) in terms of a self-adjoint (1,1)-tensor $\widetilde{Q}=-f^{2}-{I}$ reads
$T(\widetilde{Q}X,Y,Z)=T(X,\widetilde{Q}Y,Z)=T(X,Y,\widetilde{Q}Z).$
The following lemma is used in the proof of Theorem 4.3.
Lemma 4.1.
Let $(M^{2n},f,g)$ be a weak almost Hermitian manifold, considered as a nonsymmetric pseudo-Riemannian manifold $(M,G=g+F)$, where $F(X,Y)=g(X,fY)$. If an Einstein connection $\nabla$ on $M$ satisfies the $f^{2}$-torsion condition (3.3), then
$T(fY,fZ,({I}{-}f^{2})X)=$ $\,T(Y,Z,({I}{-}f^{2})f^{2}X){-}2(\nabla^{g}_{f^{2}X}\,F)(Y,fZ){+}2(\nabla^{g}_{f^{2}X}\,F)(fY,Z)$
$+dF(Y,Z,\,({I}-f^{2})f^{2}X)-dF(fY,fZ,\,({I}-f^{2})X),$ (4.1)
$T(fY,Z,\,({I}-f^{2})fX)=$ $\,-T(Y,fZ,\,({I}-f^{2})fX)-T(Y,Z,\,({I}-f^{4})X)$
$+dF(Y,Z,\,({I}-f^{4})X)+2\,(\nabla^{g}_{({I}-f^{2})X}\,F)(Y,Z)$
$-2\,(\nabla^{g}_{f^{2}X}\,F)(fY,Z)+2\,(\nabla^{g}_{f^{2}X}\,F)(Y,fZ).$ (4.2)
Proof.
Using $Z\to fZ$ in (2.7) and $\widetilde{Q}=-f^{2}-{I}$, we get
$2(\nabla^{g}_{X}\,F)(Y,fZ)=$ $-T(fZ,X,Y)-T(X,Y,fZ)+T(Z,X,fY)+T(\widetilde{Q}Z,X,fY)$
$+T(X,fY,Z)+T(X,fY,\widetilde{Q}Z)-T(Y,Z,fX)-T(Y,\widetilde{Q}Z,fX)$
$+T(fY,fZ,fX).$
Using $Y\to fY$ in (2.7), we get
$2(\nabla^{g}_{X}\,F)(fY,Z)=$ $-T(Z,X,fY)-T(X,fY,Z)+T(fZ,X,Y)+T(fZ,X,\widetilde{Q}Y)$
$+T(X,Y,fZ)+T(X,\widetilde{Q}Y,fZ)-T(Y,Z,fX)-T(\widetilde{Q}Y,Z,fX)$
$+T(fY,fZ,fX).$
Subtracting the above equations and using the $f^{2}$-torsion condition (3.3), we get
$2\{T(Z,X,fY)+T(X,fY,Z)\}+\{T(Z,\widetilde{Q}X,fY)+T(\widetilde{Q}X,fY,Z)\}$
$-2\{T(X,Y,fZ)+T(fZ,X,Y)\}-\{T(\widetilde{Q}X,Y,fZ)+T(fZ,\widetilde{Q}X,Y)\}$
$=2(\nabla^{g}_{X}\,F)(Y,fZ)-2(\nabla^{g}_{X}\,F)(fY,Z).$
Applying (2.8) to the above equation and using the $f^{2}$-torsion condition (3.3), we get
$T(fY,Z,\,({I}-f^{2})X)=$ $\,T(Y,fZ,\,({I}-f^{2})X)-2(\nabla^{g}_{X}\,F)(Y,fZ)+2(\nabla^{g}_{X}\,F)(fY,Z)$
$+dF(Y,fZ,\,({I}-f^{2})X)-dF(fY,Z,\,({I}-f^{2})X).$ (4.3)
Applying $Z\to fZ$ to (4) gives (4.1). Using (2.8) and (4.1), we rewrite (2.7) as (4.1). ∎
Remark 4.2.
For $f=J$, (4.1) and (4.1) reduce to the following:
$T(JY,JZ,X)=$ $\,-T(Y,Z,X)-dF(Y,Z,X)-dF(JY,JZ,X),$
$T(JY,Z,JX)=$ $\,-T(Y,JZ,JX){+}2\,(\nabla^{g}_{X}\,F)(Y,Z){+}(\nabla^{g}_{X}\,F)(JY,Z){-}(\nabla^{g}_{X}\,F)(Y,JZ).$
From (4.1) with $f=J$ we get
$T(JY,JZ,X)=$ $\,-T(Y,Z,X)-dF(Y,Z,X)-dF(JY,JZ,X)$
$+(\nabla^{g}_{X}\,F)(Y,JZ)-(\nabla^{g}_{X}\,F)(JY,Z).$
Subtracting the above equation from (3.19), we get a compact presentation of (2.8):
$2\,T(Y,Z,X)=(\nabla^{g}_{2\,JX-X}\,F)(JY,Z)+(\nabla^{g}_{X}\,F)(Y,JZ)-dF(Y,Z,X)-dF(JY,JZ,X).$
The following theorem generalizes Theorem 2.8, it is valid for arbitrary $F$ satisfying (3.3), and for a non-degenerate tensor $P={I}-f^{2}$ we can use it to completely determine $T$.
Theorem 4.3.
Let $\nabla$ be an Einstein connection of a nonsymmetric pseudo-Riemannian space $(M,G=g+F)$, satisfying the $f^{2}$-torsion condition (3.3), where $g(X,fY)=F(X,Y)$, and let the tensor $P:={I}-f^{2}$ have rank $2n$. If $fX\neq 0$, then the torsion $T$ of $\nabla$ is given by
$2\,T(Y,Z,\,f^{6}X)=-(\nabla^{g}_{X}F)((f+f^{3})Y,fZ)-(\nabla^{g}_{Y}F)(f^{2}Z,X)-(\nabla^{g}_{Z}F)(f^{2}X,Y)$
$+(\nabla^{g}_{fX}F)(f^{3}Y,Z)+(\nabla^{g}_{fX}F)(f^{2}Y,fZ)-(\nabla^{g}_{fY}F)(f^{3}Z,X)-(\nabla^{g}_{fZ}F)(f^{2}X,fY)$
$-(\nabla^{g}F)(P^{-1}(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X,fY,Z)+(\nabla^{g}F)(P^{-1}(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X,Y,fZ)$
$-(\nabla^{g}F)(P^{-1}\widetilde{Q}f^{4}X,fY,Z)+(\nabla^{g}F)(P^{-1}\widetilde{Q}f^{4}X,Y,fZ)+(\nabla^{g}F)(\widetilde{Q}f^{2}X,Y,Z)$
$-\tfrac{1}{2}\,dF(Y,Z,(4\,\widetilde{Q}+3\,\widetilde{Q}^{2})f^{2}X)+\tfrac{1}{2}\,dF(Y,fZ,\widetilde{Q}f^{3}X)+\tfrac{1}{2}\,dF(fY,Z,\widetilde{Q}f^{3}X)$
$+dF(fY,fZ,(\widetilde{Q}+\widetilde{Q}^{2})X),$ (4.4)
where $\widetilde{Q}=-f^{2}-{I}$, and if $fX=0$, then
$T(Y,Z,X)=2(\nabla^{g}_{X}\,F)(Y,Z)+dF(Y,Z,X).$ (4.5)
Proof.
For $fX\neq 0$, it is carried out in a similar, but more computationally complex way than the proof of Theorem 2.8. Applying $Y\to fY$ and $Z\to fZ$ to (2.7), we have
$2(\nabla^{g}_{X}F)(fY,fZ)=-T(Z,X,Y)-T(X,Y,Z)$
$\qquad-T(fZ,X,fY)-T(X,fY,fZ)-T(fY,Z,fX)-T(Y,fZ,fX)$
$\qquad-T(\widetilde{Q}Z,X,Y)-T(Z,X,\widetilde{Q}Y)-T(X,\widetilde{Q}Y,Z)-T(X,Y,\widetilde{Q}Z)$
$\qquad-T(\widetilde{Q}Z,X,\widetilde{Q}Y)-T(X,\widetilde{Q}Y,\widetilde{Q}Z)-T(fY,\widetilde{Q}Z,fX)-T(\widetilde{Q}Y,fZ,fX).$ (4.6)
From (2.7) with $Y\to-f^{2}Y$, we get
$-2(\nabla^{g}_{X}\,F)(f^{2}Y,Z)=-T(Z,X,Y)-T(X,Y,Z)-T(fZ,X,fY)-T(X,fY,fZ)$
$\qquad+T(Y,fZ,fX)+T(fY,Z,fX)-T(Z,X,\widetilde{Q}Y)-T(X,\widetilde{Q}Y,Z)$
$\qquad-T(fZ,X,\widetilde{Q}fY)-T(X,\widetilde{Q}fY,fZ)+T(\widetilde{Q}Y,fZ,fX)+T(\widetilde{Q}fY,Z,fX).$ (4.7)
The relation (4) together with (4) yields the following:
$-T(X,fY,fZ)-T(fZ,X,fY)=T(X,Y,Z)+T(Z,X,Y)+T(Z,X,\widetilde{Q}Y)$
$+T(X,\widetilde{Q}Y,Z)+\tfrac{1}{2}\big\{T(fZ,X,\widetilde{Q}fY)+T(X,\widetilde{Q}fY,fZ)+T(\widetilde{Q}Z,X,Y)$
$+T(\widetilde{Q}Z,X,\widetilde{Q}Y)+T(X,Y,\widetilde{Q}Z)+T(X,\widetilde{Q}Y,\widetilde{Q}Z)-T(\widetilde{Q}fY,Z,fX)$
$+T(fY,\widetilde{Q}Z,fX)\big\}+(\nabla^{g}_{X}F)(fY,fZ)-(\nabla^{g}_{X}\,F)(f^{2}Y,Z).$ (4.8)
Using (4), we reduce (4) to the following:
$-(\nabla^{g}_{X}\,F)(f^{2}Y,Z){-}(\nabla^{g}_{X}\,F)(fY,fZ)=T(Y,fZ,fX){+}T(fY,Z,fX){+}T(\widetilde{Q}Y,fZ,fX)$
$\quad+\tfrac{1}{2}\big\{T(\widetilde{Q}Z,X,\widetilde{Q}Y)-T(fZ,X,\widetilde{Q}fY)-T(X,\widetilde{Q}fY,fZ)+T(\widetilde{Q}Z,X,Y)$
$\quad+T(X,Y,\widetilde{Q}Z)+T(X,\widetilde{Q}Y,\widetilde{Q}Z)+T(\widetilde{Q}fY,Z,fX)+T(fY,\widetilde{Q}Z,fX)\big\}.$ (4.9)
Taking the cyclic sum of (4), in view of (4), we get:
$-(\nabla^{g}_{X}\,F)(f^{2}Y,Z)-(\nabla^{g}_{Y}\,F)(f^{2}Z,X)-(\nabla^{g}_{Z}\,F)(f^{2}X,Y)$
$\qquad=-T(Y,Z,X)-T(X,Y,Z)-T(Z,X,Y)+\tfrac{1}{2}\,\delta_{1}(X,Y,Z),$ (4.10)
where $\delta_{1}(X,Y,Z)$ is given in Section 5.
From (4) with $Y\to fY$ and $Z\to fZ$, using (4), we obtain
$-(\nabla^{g}_{X}\,F)(f^{3}Y,fZ)-(\nabla^{g}_{fY}\,F)(f^{3}Z,X)-(\nabla^{g}_{fZ}\,F)(f^{2}X,fY)$
$+(\nabla^{g}_{X}\,F)(f^{2}Y,Z)-(\nabla^{g}_{X}\,F)(fY,fZ)$
$\quad=T(X,Y,Z)+T(Z,X,Y)-T(fY,fZ,X)+\tfrac{1}{2}\,\delta_{2}(X,Y,Z),$ (4.11)
where $\delta_{2}(X,Y,Z)$ is given in Section 5. Replacing $X\to fX$ and $Y\to fY$ in (4), we get:
$T(fY,fZ,X)=T(Y,Z,X)+(\nabla^{g}_{fX}\,F)(f^{2}Y,fZ)+(\nabla^{g}_{fX}\,F)(f^{3}Y,Z)$
$+\tfrac{1}{2}\,T(fZ,fX,\widetilde{Q}Y)+\tfrac{1}{2}\,T(fZ,fX,\widetilde{Q}^{2}Y)+\tfrac{1}{2}\,T(fX,\widetilde{Q}Y,fZ)+\tfrac{1}{2}\,T(fX,\widetilde{Q}^{2}Y,fZ)$
$+\tfrac{3}{2}\,T(\widetilde{Q}Y,Z,X)+\tfrac{1}{2}\,T(\widetilde{Q}^{2}Y,Z,X)+\tfrac{3}{2}\,T(\widetilde{Q}Y,Z,\widetilde{Q}X)+\tfrac{1}{2}\,T(\widetilde{Q}^{2}Y,Z,\widetilde{Q}X)$
$+\tfrac{1}{2}\,T(\widetilde{Q}Z,fX,fY)+\tfrac{1}{2}\,T(\widetilde{Q}Z,fX,\widetilde{Q}fY)+\tfrac{1}{2}\,T(fX,fY,\widetilde{Q}Z)+\tfrac{1}{2}\,T(fX,\widetilde{Q}fY,\widetilde{Q}Z)$
$+\tfrac{1}{2}\,T(Y,\widetilde{Q}Z,X)+\tfrac{1}{2}\,T(\widetilde{Q}Y,\widetilde{Q}Z,X)+\tfrac{1}{2}\,T(Y,\widetilde{Q}Z,\widetilde{Q}X)+\tfrac{1}{2}\,T(\widetilde{Q}Y,\widetilde{Q}Z,\widetilde{Q}X)$
$-T(fY,fZ,\widetilde{Q}X)-T(\widetilde{Q}fY,fZ,X)-T(\widetilde{Q}fY,fZ,\widetilde{Q}X)+T(Y,Z,\widetilde{Q}X).$ (4.12)
Then substituting (4) into (4), we get the following:
$(\nabla^{g}_{X}F)(f^{2}Y,Z)-(\nabla^{g}_{X}F)((f+f^{3})Y,fZ)-(\nabla^{g}_{fY}F)(f^{3}Z,X)$
$-(\nabla^{g}_{fZ}F)(f^{2}X,fY)+(\nabla^{g}_{fX}F)(f^{3}Y,Z)+(\nabla^{g}_{fX}F)(f^{2}Y,fZ)$
$\quad=T(X,Y,Z)+T(Z,X,Y)-T(Y,Z,X)+\tfrac{1}{2}\,\delta_{3}(X,Y,Z),$
where $\delta_{3}(X,Y,Z)$ is given in Section 5. This relation together with (4) yields
$2\,T(Y,Z,X)=$ $\ (\nabla^{g}_{X}F)((f+f^{3})Y,fZ)+(\nabla^{g}_{Y}F)(f^{2}Z,X)+(\nabla^{g}_{Z}F)(f^{2}X,Y)$
$-(\nabla^{g}_{fX}F)(f^{3}Y,Z)-(\nabla^{g}_{fX}F)(f^{2}Y,fZ)+(\nabla^{g}_{fY}F)(f^{3}Z,X)$
$+(\nabla^{g}_{fZ}F)(f^{2}X,fY)+\tfrac{1}{2}\,\delta_{4}(X,Y,Z),$ (4.13)
where $\delta_{4}(X,Y,Z)$ is given in Section 5. Using (2.8) and the $f^{2}$-torsion condition, we reduce $\delta_{4}(X,Y,Z)$ to $\delta_{5}(X,Y,Z)$, given in ection 5. Using (4.1) and (4.1), we find
$T(fY,fZ,(3\,\widetilde{Q}$ $+2\,\widetilde{Q}^{2})X)=T(Y,Z,(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X)+dF(Y,Z,(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X)$
$-dF(fY,fZ,(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})X)-2\,(\nabla^{g}F)(P^{-1}(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X,Y,fZ)$
$+2\,(\nabla^{g}F)(P^{-1}(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X,fY,Z),$
$T(Y,fZ,\widetilde{Q}f^{3}X)$ $+T(fY,Z,\widetilde{Q}f^{3}X)=dF(Y,Z,\widetilde{Q}f^{2}({I}+f^{2})X)+2\,(\nabla^{g}F)(\widetilde{Q}f^{2}X,Y,Z)$
$+2\,(\nabla^{g}F)(P^{-1}\widetilde{Q}f^{4}X,Y,fZ)-2\,(\nabla^{g}F)(P^{-1}\widetilde{Q}f^{4}X,fY,Z),$
and then rewrite $\delta_{5}(X,Y,Z)$ as
$\delta_{5}(X,Y,Z)=$ $\,-T(Y,Z,(9\,\widetilde{Q}+7\,\widetilde{Q}^{2}+2\,\widetilde{Q}^{3})X)+T(Y,Z,(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X)$
$+dF(Y,Z,(3\,\widetilde{Q}{+}2\,\widetilde{Q}^{2})f^{2}X){-}dF(Y,Z,\widetilde{Q}f^{2}({I}+f^{2})X){+}dF(Y,Z,\widetilde{Q}f^{2}X)$
$-2\,dF(fY,fZ,(\widetilde{Q}+\widetilde{Q}^{2})X)-2\,(\nabla^{g}F)(P^{-1}(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X,Y,fZ)$
$+2\,(\nabla^{g}F)(P^{-1}(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X,fY,Z)-2\,(\nabla^{g}F)(\widetilde{Q}f^{2}X,Y,Z)$
$-2\,(\nabla^{g}F)(P^{-1}\widetilde{Q}f^{4}X,Y,fZ)+2\,(\nabla^{g}F)(P^{-1}\widetilde{Q}f^{4}X,fY,Z)$
$-dF(Y,fZ,\widetilde{Q}f^{3}X)-dF(fY,Z,\widetilde{Q}f^{3}X).$
Note that $\delta_{i}(X,Y,Z)\ (1\leq i\leq 5)$ vanish when $\widetilde{Q}=0$. From the equation above and (4) with $\delta_{4}(X,Y,Z)$ replaced by $\delta_{5}(X,Y,Z)$, the equation (4.3) follows. If $fX=0$, then from (4.1) we directly get (4.5). ∎
Example 4.4.
Take almost Hermitian manifolds $(M_{j},J_{j},g_{j})\ (j=1,\dots,m)$, where $J_{j}^{2}=-{I}_{\,j}$. Consider the Riemannian product $(M,g)=\bigl(M_{1}\times\cdots\times M_{m},\;g_{1}\oplus\cdots\oplus g_{m}\bigr)$, and define the (1,1)(1,1)–tensor $f=\sqrt{\lambda_{1}}\,J_{1}\oplus\cdots\oplus\sqrt{\lambda_{m}}\,J_{m}$. Then $(M,f,g)$ is a weak almost Hermitian manifold with $f^{2}=-\bigoplus_{j}\lambda_{j}\,{I}_{j}$, and on $TM_{j}$ (and for $X\in TM_{j}$) we have
$fX=\sqrt{\lambda_{j}}\,J_{j}X,\quad f^{2}X=-\lambda_{j}X,\quad f^{3}X=-\lambda_{j}^{3/2}J_{j}X,\quad f^{4}X=\lambda_{j}^{2}X,\quad f^{6}X=-\lambda_{j}^{3}X,$
$\widetilde{Q}|_{\,TM_{j}}=(\lambda_{j}-1)\,{I}_{j},\quad P={I}-f^{2}=\bigoplus\nolimits_{j}(1+\lambda_{j})\,{I}_{j},\quad P^{-1}=\bigoplus\nolimits_{j}\frac{1}{1+\lambda_{j}}\,{I}_{j}.$
Since the Levi-Civita connection of the Riemannian product preserves the splitting $TM=\bigoplus_{j=1}^{m}TM_{j}$, $f^{2}$ acts as a multiple of the identity on each factor. The 2-form $F$ corresponding to $f$ is given by $F=F_{1}\oplus\cdots\oplus F_{m}$, where $F_{j}(X,Y)=g_{j}(X,\sqrt{\lambda}_{j}J_{j}Y)$. As a consequence,
$(\nabla^{g}_{X}F)(Y,Z)=0,$
whenever $X$ belongs to one factor and at least one of $Y,Z$ belongs to a different factor.
If an Einstein connection $\nabla$ satisfies the $f^{2}$-torsion condition (3.3), then $T(X,Y)=0$ for $X\in TM_{i},\,Y\in TM_{j}\ (i\neq j)$. Hence, the torsion tensor splits into the direct sum $T=T|_{TM_{1}}\oplus\cdots\oplus T|_{TM_{m}}$. Substituting $\widetilde{Q}|_{\,TM_{j}}=(\lambda_{j}-1)\,{I}_{j}$ into (4.3), then using
$(\nabla^{g_{j}}_{X}F_{j})(J_{j}Y,J_{j}Z)=-\,(\nabla^{g_{j}}_{X}F_{j})(Y,Z),\quad(\nabla^{g_{j}}_{X}F_{j})(J_{j}Y,Z)=(\nabla^{g_{j}}_{X}F_{j})(Y,J_{j}Z),$
we uniquely determine the torsion component $T|_{TM_{j}}$ by
$T|_{\,TM_{j}}(Y,Z,X)$ $=\lambda_{j}^{-2}(\lambda_{j}-1)(\nabla^{g_{j}}_{X}F_{j})(Y,Z)-\tfrac{1}{2}\lambda_{j}^{-2}(\nabla^{g_{j}}_{Y}F_{j})(Z,X)-\tfrac{1}{2}\lambda_{j}^{-2}(\nabla^{g_{j}}_{Z}F_{j})(X,Y)$
$\ +\lambda_{j}^{-1}(\nabla^{g_{j}}_{J_{j}X}F_{j})(Y,J_{j}Z){-}\tfrac{1}{2}\lambda_{j}^{-1}(\nabla^{g_{j}}_{J_{j}Y}F_{j})(J_{j}Z,X){-}\tfrac{1}{2}\lambda_{j}^{-1}(\nabla^{g_{j}}_{J_{j}Z}F_{j})(X,J_{j}Y)$
$\ -\tfrac{1}{4}\,\lambda_{j}^{-2}(\lambda_{j}-1)(3\lambda_{j}+1)dF_{j}(Y,Z,X)+\tfrac{1}{4}(\lambda_{j}-1)\lambda_{j}^{-1/2}dF_{j}(Y,J_{j}Z,J_{j}X)$
$\ +\tfrac{1}{4}(\lambda_{j}-1)\lambda_{j}^{-1/2}dF_{j}(J_{j}Y,Z,J_{j}X)-\tfrac{1}{2}\lambda_{j}^{-2}(\lambda_{j}-1)dF_{j}(J_{j}Y,J_{j}Z,X).$
If each $(M_{j},J_{j},g_{j})$ is a Kähler manifold, then $\nabla^{g_{j}}F_{j}=0$ and $dF_{j}=0$. In this case, all torsion components vanish, and the Einstein connection reduces to the Levi-Civita connection. Using (2.5), we can further simplify the formula for $T|_{TM_{j}}$ as
$T|_{TM_{j}}(Y,Z,X)=$ $\ \frac{(\lambda-1)(5-3\lambda)}{4\lambda^{2}}\,(\nabla^{g_{j}}_{X}F_{j})(Y,Z)+\Bigl(\frac{1}{\lambda}+\frac{\lambda-1}{2\sqrt{\lambda}}\Bigr)\,(\nabla^{g_{j}}_{J_{j}X}F_{j})(Y,J_{j}Z)$
$-\Bigl(\frac{3\lambda^{2}-2\lambda+1}{4\lambda^{2}}+\frac{\lambda-1}{4\sqrt{\lambda}}\Bigr)\,\big\{(\nabla^{g_{j}}_{Z}F_{j})(X,Y)+(\nabla^{g_{j}}_{Y}F_{j})(Z,X)\big\}$
$+\Bigl(\frac{\lambda-1}{4\sqrt{\lambda}}{-}\frac{1}{2\lambda}{-}\frac{\lambda-1}{2\lambda^{2}}\Bigr)\,\big\{(\nabla^{g_{j}}_{J_{j}Y}F_{j})(J_{j}Z,X)+(\nabla^{g_{j}}_{J_{j}Z}F_{j})(X,J_{j}Y)\big\}.$
The formula above for $\lambda=1$ gives the solution (2.8).
Corollary 4.5.
Let $\nabla$ be an Einstein connection of a weak almost Hermitian manifold $(M^{2n},f,g>0)$, considered as a nonsymmetric Riemannian space $(M,G=g+F)$ with $F(X,Y)=g(X,fY)$. If the condition (3.3) is valid, then the torsion $T$ of $\nabla$ is given by (4.3).
Remark 4.6.
For a nonsymmetric Riemannian space $(M,G=g+F)$, the tensor $P={I}-f^{2}$ is positive definite, and therefore non-degenerate. Using (2.5), we can further simplify (4.3) to the form with $\nabla^{g}\,F$-terms.
Corollary 4.7 (see [12]).
Let $(M^{2n},f,g)$ be a weak almost Hermitian manifold with fundamental $2$-form $F$, considered as an NGT space $(M,G=g+F,\nabla)$, and $\nabla$ (an Einstein connection) has a totally skew-symmetric torsion $(0,3)$-tensor $T$. If the $f$-torsion condition (2.17) is true, then $(M^{2n},f,g)$ is a weak nearly Kähler manifold and the following is true:
$T(fX,Y,Z)=-\tfrac{1}{3}\,dF(fX,Y,Z),\quad\nabla\,f^{2}=\nabla^{g}f^{2}=0.$
5 The terms $\delta_{i}(X,Y,Z)$
Here, we present the terms $\delta_{i}(X,Y,Z)\ (1\leq i\leq 5)$, see Section 4. Their expressions were verified with the help of MAPLE program.
$\delta_{1}(X,Y,Z)=$ $\,-T(fZ,X,\widetilde{Q}fY)-T(X,\widetilde{Q}Y,Z)-T(X,\widetilde{Q}fY,fZ)+T(\widetilde{Q}fY,Z,fX)$
$-T(Z,X,\widetilde{Q}Y)-T(X,Y,\widetilde{Q}Z)+T(\widetilde{Q}Y,fZ,fX)-T(fX,Y,\widetilde{Q}fZ)$
$-T(Y,\widetilde{Q}Z,X)-T(Y,\widetilde{Q}fZ,fX)-T(Y,Z,\widetilde{Q}X)+T(\widetilde{Q}fZ,X,fY)$
$-T(Z,\widetilde{Q}X,Y)+T(\widetilde{Q}fX,Y,fZ)-T(fY,Z,\widetilde{Q}fX)-T(Z,\widetilde{Q}fX,fY)$
$+T(\widetilde{Q}Z,fX,fY)+T(\widetilde{Q}X,fY,fZ).$
$\delta_{2}(X,Y,Z)=$ $\ T(X,\widetilde{Q}Y,Z)-T(\widetilde{Q}Y,fZ,fX)-2\,T(\widetilde{Q}fY,Z,fX)+T(Z,X,\widetilde{Q}Y)$
$+2\,T(\widetilde{Q}Z,X,Y){+}T(\widetilde{Q}Z,X,\widetilde{Q}Y){+}T(X,Y,\widetilde{Q}Z){+}2\,T(fY,\widetilde{Q}Z,fX)$
$-T(\widetilde{Q}fX,fY,Z)-T(\widetilde{Q}fX,fY,\widetilde{Q}Z)+T(fX,fY,\widetilde{Q}Z)$
$+T(fX,fY,\widetilde{Q}^{2}Z)+T(fY,\widetilde{Q}^{2}Z,fX)+T(Y,fZ,\widetilde{Q}fX)$
$+T(\widetilde{Q}Y,fZ,\widetilde{Q}fX)-T(\widetilde{Q}^{2}Y,fZ,fX)+T(fZ,\widetilde{Q}fX,Y)$
$+T(fZ,\widetilde{Q}fX,\widetilde{Q}Y)-T(\widetilde{Q}fY,\widetilde{Q}Z,fX)-T(\widetilde{Q}fZ,fX,Y)$
$-T(\widetilde{Q}fZ,fX,\widetilde{Q}Y)-T(X,fY,\widetilde{Q}fZ)-T(fY,\widetilde{Q}fZ,X)$
$-T(fY,fZ,\widetilde{Q}X)+T(\widetilde{Q}^{2}Z,X,Y)+T(\widetilde{Q}^{2}Z,X,\widetilde{Q}Y)$
$-T(X,\widetilde{Q}^{2}Y,Z)-T(X,\widetilde{Q}^{2}Y,\widetilde{Q}Z)+T(\widetilde{Q}X,\widetilde{Q}Y,Z)$
$+T(\widetilde{Q}X,\widetilde{Q}Y,\widetilde{Q}Z)-T(Z,X,\widetilde{Q}^{2}Y)-T(\widetilde{Q}Z,X,\widetilde{Q}^{2}Y)$
$+T(\widetilde{Q}X,Y,Z)+T(\widetilde{Q}X,Y,\widetilde{Q}Z)-T(fZ,\widetilde{Q}X,fY),$
$\delta_{3}(X,Y,Z)=$ $\ T(fZ,\widetilde{Q}fX,\widetilde{Q}Y){-}T(\widetilde{Q}fZ,fX,\widetilde{Q}Y){-}T(\widetilde{Q}fZ,fX,Y){-}T(\widetilde{Q}fY,\widetilde{Q}Z,fX)$
$+T(fZ,\widetilde{Q}fX,Y){-}T(\widetilde{Q}^{2}Y,fZ,fX){+}T(\widetilde{Q}Y,fZ,\widetilde{Q}fX){+}T(Y,fZ,\widetilde{Q}fX)$
$+T(fY,\widetilde{Q}^{2}Z,fX){+}T(fX,fY,\widetilde{Q}^{2}Z){-}T(\widetilde{Q}fX,fY,\widetilde{Q}Z){-}T(\widetilde{Q}fX,fY,Z)$
$-T(\widetilde{Q}Z,X,\widetilde{Q}^{2}Y)-T(Z,X,\widetilde{Q}^{2}Y)-3\,T(\widetilde{Q}Y,Z,\widetilde{Q}X)+T(\widetilde{Q}X,\widetilde{Q}Y,\widetilde{Q}Z)$
$-T(fX,\widetilde{Q}Y,fZ)-3\,T(\widetilde{Q}Y,Z,X)-T(fX,\widetilde{Q}fY,\widetilde{Q}Z)-T(\widetilde{Q}Z,fX,\widetilde{Q}fY)$
$+2\,T(\widetilde{Q}fY,fZ,\widetilde{Q}X){+}2\,T(\widetilde{Q}fY,fZ,X){-}T(fX,\widetilde{Q}^{2}Y,fZ){-}T(fZ,fX,\widetilde{Q}Y)$
$-T(fZ,fX,\widetilde{Q}^{2}Y)-T(\widetilde{Q}^{2}Y,Z,X)-T(\widetilde{Q}^{2}Y,Z,\widetilde{Q}X)-T(\widetilde{Q}Y,\widetilde{Q}Z,X)$
$-T(\widetilde{Q}Y,\widetilde{Q}Z,\widetilde{Q}X)-T(Y,\widetilde{Q}Z,\widetilde{Q}X)+T(\widetilde{Q}X,\widetilde{Q}Y,Z)-T(X,\widetilde{Q}^{2}Y,\widetilde{Q}Z)$
$-T(X,\widetilde{Q}^{2}Y,Z)+T(\widetilde{Q}^{2}Z,X,\widetilde{Q}Y)+T(\widetilde{Q}^{2}Z,X,Y)+T(fY,fZ,\widetilde{Q}X)$
$-T(fY,\widetilde{Q}fZ,X)-T(X,fY,\widetilde{Q}fZ)-T(fZ,\widetilde{Q}X,fY)+T(\widetilde{Q}X,Y,\widetilde{Q}Z)$
$+T(\widetilde{Q}X,Y,Z)-T(\widetilde{Q}Y,fZ,fX)+T(X,Y,\widetilde{Q}Z)+T(\widetilde{Q}Z,X,\widetilde{Q}Y)$
$+T(Z,X,\widetilde{Q}Y)+T(X,\widetilde{Q}Y,Z)+2\,T(fY,\widetilde{Q}Z,fX)+2\,T(\widetilde{Q}Z,X,Y)$
$-2\,T(\widetilde{Q}fY,Z,fX)-T(\widetilde{Q}Z,fX,fY)-2\,T(Y,Z,\widetilde{Q}X)-T(Y,\widetilde{Q}Z,X),$
$\delta_{4}(X,Y,Z)=$ $\ 2\,T(\widetilde{Q}Z,X,Y)-T(fZ,X,\widetilde{Q}fY)-T(X,\widetilde{Q}fY,fZ)-T(\widetilde{Q}fY,Z,fX)$
$+T(\widetilde{Q}Z,X,\widetilde{Q}Y)+2\,T(fY,\widetilde{Q}Z,fX)-T(fX,Y,\widetilde{Q}fZ)-2\,T(Y,\widetilde{Q}Z,X)$
$-T(Y,\widetilde{Q}fZ,fX)+T(\widetilde{Q}X,Y,Z)+T(\widetilde{Q}X,Y,\widetilde{Q}Z)-3\,T(Y,Z,\widetilde{Q}X)$
$-T(Y,\widetilde{Q}Z,\widetilde{Q}X)+T(\widetilde{Q}fZ,X,fY)-T(fZ,\widetilde{Q}X,fY)-T(Z,\widetilde{Q}X,Y)$
$+T(\widetilde{Q}fX,Y,fZ)-T(fY,Z,\widetilde{Q}fX)-T(Z,\widetilde{Q}fX,fY)-3\,T(\widetilde{Q}Y,Z,X)$
$-3\,T(\widetilde{Q}Y,Z,\widetilde{Q}X)-T(fX,\widetilde{Q}Y,fZ)+T(\widetilde{Q}X,fY,fZ)-T(X,fY,\widetilde{Q}fZ)$
$-T(fY,\widetilde{Q}fZ,X)+T(fY,fZ,\widetilde{Q}X)+T(\widetilde{Q}^{2}Z,X,Y)+T(\widetilde{Q}^{2}Z,X,\widetilde{Q}Y)$
$-T(X,\widetilde{Q}^{2}Y,Z)-T(X,\widetilde{Q}^{2}Y,\widetilde{Q}Z)+T(\widetilde{Q}X,\widetilde{Q}Y,Z)+T(\widetilde{Q}X,\widetilde{Q}Y,\widetilde{Q}Z)$
$-T(Z,X,\widetilde{Q}^{2}Y)-T(\widetilde{Q}Z,X,\widetilde{Q}^{2}Y)-T(\widetilde{Q}fX,fY,Z)-T(\widetilde{Q}fX,fY,\widetilde{Q}Z)$
$+T(fX,fY,\widetilde{Q}^{2}Z){+}T(fY,\widetilde{Q}^{2}Z,fX){+}T(Y,fZ,\widetilde{Q}fX){+}T(\widetilde{Q}Y,fZ,\widetilde{Q}fX)$
$-T(\widetilde{Q}^{2}Y,fZ,fX){+}T(fZ,\widetilde{Q}fX,Y){+}T(fZ,\widetilde{Q}fX,\widetilde{Q}Y){-}T(fX,\widetilde{Q}^{2}Y,fZ)$
$-T(fZ,fX,\widetilde{Q}Y)-T(fZ,fX,\widetilde{Q}^{2}Y)-T(\widetilde{Q}^{2}Y,Z,X)-T(\widetilde{Q}^{2}Y,Z,\widetilde{Q}X)$
$-T(\widetilde{Q}Y,\widetilde{Q}Z,X)-T(\widetilde{Q}Y,\widetilde{Q}Z,\widetilde{Q}X)+2\,T(\widetilde{Q}fY,fZ,X)$
$+2\,T(\widetilde{Q}fY,fZ,\widetilde{Q}X)-T(\widetilde{Q}fY,\widetilde{Q}Z,fX)-T(\widetilde{Q}fZ,fX,Y)$
$-T(\widetilde{Q}fZ,fX,\widetilde{Q}Y)-T(\widetilde{Q}Z,fX,\widetilde{Q}fY)-T(fX,\widetilde{Q}fY,\widetilde{Q}Z),$
[…]
$-T(fX,Y,\widetilde{Q}fZ)-T(X,fY,\widetilde{Q}fZ)+T(Z,X,\widetilde{Q}^{2}Y)-T(fX,fY,\widetilde{Q}Z)$
$-T(fX,fY,\widetilde{Q}^{2}Z){+}2\,T(fY,fZ,\widetilde{Q}X){-}T(fZ,fX,\widetilde{Q}Y){-}T(fZ,fX,\widetilde{Q}^{2}Y)$
$+2\,T(fY,fZ,\widetilde{Q}^{2}X){-}6\,T(Y,Z,\widetilde{Q}^{2}X){-}2\,T(Y,Z,\widetilde{Q}^{3}X){-}T(Z,fX,\widetilde{Q}^{2}fY)$
$+T(X,Y,\widetilde{Q}^{2}Z)-T(fX,Y,\widetilde{Q}^{2}fZ)-T(Z,fX,\widetilde{Q}fY)$
[…]
$-T(Y,fZ,\widetilde{Q}f^{3}X)-T(fY,Z,\widetilde{Q}f^{3}X)-dF(Y,fZ,\widetilde{Q}f^{3}X)$
$-dF(fY,Z,\widetilde{Q}f^{3}X)+dF(fY,fZ,\widetilde{Q}X)+dF(Y,Z,\widetilde{Q}f^{2}X).$
Conclusion
We presented explicitly the Einstein connection for nonsymmetric pseudo-Riemannian, in particular, for weak almost Hermitian manifolds, satisfying the $f^{2}$-torsion condition. We also obtained the corresponding formulas for almost contact metric manifolds, assuming the $f^{2}$-torsion condition. We expressed the torsion in terms of $\nabla^{g}F$ and $dF$, and the additional $\widetilde{Q}-$terms, explained precisely how the weak almost Hermitian case differs from the almost Hermitian case. The presented identities provide a new tool for constructing examples and for studying classes of nonsymmetric pseudo-Riemannian manifolds, in particular, weak almost Hermitian manifolds with the usual Gray-Hervella classification, in Einstein’s nonsymmetric gravitational theory. In further research we will study weak (para-) metric $f$-manifolds, in particular, weak (para-) contact metric manifolds, with an Einstein connections or connections with totally skew-symmetric torsion, satisfying the $f^{2}$-torsion condition for vectors orthogonal to the distribution $\ker f$.
Acknowledgments. This work was partially supported by the Ministry of Education, Science and Technological Development of the Republic of Serbia (contract reg. no. 451-03-34/2026-03/200124 for Milan Zlatanović). The authors would also like to thank Miroslav Maksimović for carefully reading the paper and for his helpful and valuable comments.